OpenStudy (anonymous):

what is slope -2/3,yintercept(0,5)

OpenStudy (anonymous):

\[y-5=-2/3(x-0)\]

OpenStudy (anonymous):

\[y=-2/3x+5\]

OpenStudy (anonymous):

is that what you were looking for? the equation of the line?

OpenStudy (anonymous):

i can show you how to do that

OpenStudy (anonymous):

you use the "point slope" formula which is y – y1 = m(x – x1)

OpenStudy (anonymous):

first find the gradient between those line... \[m = \frac{ x _{2}-x_{1} }{ y _{2}-y_{1} }\] then by substituting one of the point to the general linear equation : \[y-y _{1} = m(x-x _{1})\] your answer should be... \[y=x+5\] *try it yourself...

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

y1 is the your y value @ the point (0, 5) and x1 is the x value at the point (0 ,5) M is the slope plug in the numbers and simplify

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

so is m always going to be the slope